# Ban&Pick: Achieving Free Performance Gains and Inference Speedup via Smarter Routing in MoE-LLMs

### v1

@@ Abstract
Sparse Mixture-of-Experts (MoE) has become a key architecture for scaling large language models (LLMs) efficiently. Recent fine‑grained MoE designs introduce hundreds of experts per layer, with multiple experts activated per token, enabling stronger specialization. However, during pre‑training, routers are optimized mainly for stability and robustness: they converge prematurely and enforce balanced usage, limiting the full potential of model performance and efficiency. In this work, we uncover two overlooked issues: (i) a few highly influential experts are underutilized due to premature and balanced routing decisions; and (ii) enforcing a fixed number of active experts per token introduces substantial redundancy. Instead of retraining models or redesigning MoE architectures, we introduce Ban&Pick, a post-training, plug-and-play strategy for smarter MoE routing. Pick discovers and reinforces key experts—a small group with outsized impact on performance—leading to notable accuracy gains across domains. Ban complements this by dynamically pruning redundant experts based on layer and token sensitivity, delivering faster inference with minimal accuracy loss. Experiments on fine-grained MoE-LLMs (DeepSeek, Qwen3) across math, code, and general reasoning benchmarks demonstrate that Ban&Pick delivers free performance gains and inference acceleration without retraining or architectural changes. For instance, on Qwen3-30B-A3B, it improves accuracy from 80.67 to 84.66 on AIME2024 and from 65.66 to 68.18 on GPQA-Diamond, while accelerating inference by 1.25× under the vLLM.
@@ 1 Introduction
Large language models have achieved remarkable progress across diverse domains (DBLP:conf/iclr/ZhouWLSLQLJSZ024, ), yet the ever-growing demand for higher capacity presents significant challenges for efficiency. As a promising solution, Mixture-of-Experts (MoE) architectures scale parameter count without proportionally increasing computation, by activating only a small subset of experts for each input token (shazeer2017outrageously, ; artetxe-etal-2022-efficient, ).
@@
Yet, whether the routing strategies learned during pre-training can fully exploit this specialization remains an open question. Prior evidence from OLMoE (muennighoff2025olmoeopenmixtureofexpertslanguage, ) and OpenMoE (xue2024openmoe, ) suggests that routers converge much earlier than the experts. In particular, OLMoE reports that after only 1% of pre-training, around 60% of the routing decisions have already saturated, while OpenMoE finds that by 40% of pre-training, about 80% of the routing has stabilized. Such early fixation implies that routing decisions are largely determined before the experts have sufficiently matured. Moreover, the widely used balancing loss (lepikhin2020gshardscalinggiantmodels, ), while effective at preventing collapse to just a few experts, may also discourage tokens from concentrating on the most impactful experts. As a result, the stronger or more specialized experts might not be fully exploited, while tokens are simultaneously spread across less relevant ones—potentially leading to redundancy in expert selection. These training-time phenomena strongly motivate us to design a smarter routing strategy that can both better leverage the most capable experts and alleviate redundancy. Instead of attempting to redesign training objectives or expert architectures (yan2025tcmoe, ; jin2024moeacceleratingmixtureofexpertsmethods, ) which often trade off train stability for flexibility—we take a post-training perspective, aiming to improve routing behaviors without retraining or modifying the architecture.
Delving into this question, we surprisingly find that within the most frequently selected experts of fine‑grained MoE models, a small fraction plays an outsized role: simply forcing these experts to be chosen for every input can noticeably boost accuracy on certain datasets. Motivated by this striking observation, we conduct an empirical analysis of expert usage across domains such as math, code, and general reasoning. We discover that each domain consistently relies on its own set of frequently selected experts, which we refer to as domain‑specialized experts. Yet, a closer examination further shows that only a tight subset of these domain‑specialized experts has a decisive impact on the final logits distribution, while the influence of the others remains marginal. We term these most critical ones key experts. Our Pick module is designed to reinforce their influence during routing, enabling the model to emphasize key experts and thereby improve domain‑specific performance.
@@
Together, we introduce Ban&Pick, a unified framework that improves both the accuracy and efficiency of MoE without retraining or architectural changes. Evaluated on five widely used datasets spanning math, coding, and general reasoning, the results demonstrate that on the Qwen3 series models, applying Pick alone brings an average 2.83% performance improvement. By further incorporating Ban, the joint Ban&Pick strategy achieves an average 1.41% improvement while simultaneously accelerating inference by 1.25×. These findings highlight that fine-grained MoE models, once routed more intelligently, can deliver stronger performance and faster inference without extra cost.
@@ 2 Related Works
Mixture-of-Experts.
The Sparsely-Gated MoE (6797059, ; 6796382, ) first introduced the idea of using a trainable gating network to activate only a small subset of experts, typically through a Top‑K routing scheme. Subsequent studies (deepseekai2024deepseekv2strongeconomicalefficient, ; yang2025qwen3technicalreport, ) built upon this framework, scaling MoE models to much larger sizes and demonstrating that sparse expert activation can deliver strong efficiency and performance gains. As these models expanded, the routing mechanism became a critical factor of overall effectiveness.
Enhancing Reasoning in MoE.
A common line of work seeks to enhance the reasoning ability of LLMs through supervised fine-tuning (SFT) on domain-specific or task-specific data (wei2022finetunedlanguagemodelszeroshot, ; muennighoff2025s1simpletesttimescaling, ). Another set of approaches improves reasoning by test‑time scaling (lee2025evolvingdeeperllmthinking, ; yao2023treethoughtsdeliberateproblem, ; Besta_2024, ; wang2025thoughtsplaceunderthinkingo1like, ), such as increasing the number of sampled responses or extending intermediate reasoning steps. Although these methods often yield stronger reasoning performance, they introduce inference overhead. More recently, (wang2025expertsneedsteeringthinking, ) proposes a method specifically designed for MoE models that employ the <think> token, enhancing MoE reasoning by amplifying the weights of Cognitive Experts.
Dynamic Experts Pruning of MoE-LLMs.
Dynamic expert pruning has emerged as one of the key techniques for accelerating MoE inference. (lu-etal-2024-experts, ) proposed a pruning strategy based on weight differences between candidate experts: simply skipping experts with less weights. Building on this, (huang2024mc, ) introduced a critical token protection mechanism. However, the substantial sensitivity differences across layers of MoE models and across tokens in long chain‑of‑thought reasoning are often overlooked, leaving room for further optimization.
3 Motivation
Figure 1: The intuition and empirical findings on expert routing. Left (Pick): forcibly activating key experts (e.g., E43) improves accuracy on Winogrande (sakaguchi2021winogrande, ). Right (Ban): reducing experts from 6 to 5 cuts 16.7% computation and delivers 10% speedup without accuracy loss.
Recent evidence suggests that the current pre‑training paradigm for MoE makes a deliberate trade‑off: in exchange for routing stability, it locks routers into decisions before experts have fully matured. Combined with balancing constraints, this often prevents tokens from concentrating on the strongest or most impactful experts, while simultaneously spreading computation across many redundant ones. In other words, stability comes at the cost of underutilizing the very experts that carry the greatest potential and of inflating the role of those that contribute little.
@@ 3 Motivation
To better understand the optimization space, we conducted two simple yet revealing experiments with the Deepseek‑V2‑Lite‑Chat model (activates 6 experts out of 64) and showed results in Figure 1.
First, we examined expert selection frequency on the C4 (dodge2021documentinglargewebtextcorpora, ) calibration set, identified the top-3 most frequently selected experts in each layer, and then conducted a grid search: each of these experts was manually added in addition to the original routing (E43 in Figure 1), and the resulting accuracy on Winogrande was recorded. The left part highlights a few experts contribute disproportionately: enhancing Layer‑1 Expert‑43 improved accuracy by +1.81%, L10E33 by +1.65%, and L25E11 by +1.26%. When these three were jointly enhanced, accuracy increased by +2.72% (71.67% $\rightarrow$ 74.39%), while most other high-frequency experts had negligible effects. This observation suggests the Pick intuition: selectively reinforcing a small subset of high-frequency experts has the potential to yield significant performance gains. We refer to this subset of experts as key experts.
Second, we reduced the number of selected experts per token from 6 to 5, and measured both the resulting accuracy and inference speedup. Remarkably, as shown in the right side of Figure 1, this pruning reduced computation by 16.7% and delivered a 10% inference speedup, all without any accuracy drop on Winogrande (still 71.67%). This observation supports the Ban intuition: dynamically pruning redundancy in expert selection can improve efficiency with minimal risk.
Together, these two experiments highlight two opportunities: (i) identifying and enhancing key experts as a potential avenue for improved performance, and (ii) pushing the limits of expert pruning to reduce cost. They motivate us to design a unified framework that can pick beneficial experts while banning redundant ones, aiming at improving both accuracy and efficiency.
@@ 4 Pick
In this section, We begin with an analysis of expert specialization in fine‑grained MoE models to examine whether domain-specialized experts indeed emerge. Building on this analysis, we then introduce our Pick method in two steps: (i) identifying key experts from the set of domain‑specialized experts, and (ii) designing strategies to enhance these key experts so that their impact on model performance can be more effectively amplified.
@@ 4.1 Expert Specialization in Fine-Grained MoE
First, we take a closer look at the degree of expert specialization in fine‑grained MoE models. Specifically, we analyze the Qwen3‑30B‑A3B model and record expert usage frequencies on three representative benchmarks: a math task (GSM8K (cobbe2021gsm8k, )), a code task (MBPP
 (austin2021program, )), and a general task (GPQA-Diamond). The statistics cover all tokens in both the prefill and decode stages. For each task, we plot the per‑layer expert usage frequency.
Our analysis reveals that in each layer there exist experts whose usage frequency is significantly higher than the average–often up to 6 times higher. Moreover, in most layers these high‑frequency experts vary across tasks. As illustrated in Figure 2, we show 3 representative layers (complete results in Appendix A). For instance, L16E95 is frequently activated only on math tasks, whereas L29E91 is almost exclusively activated on code tasks with an extremely high activation rate of 66.9%. In contrast, a small number of experts such as L4E48 remain highly active across all categories of tasks.
@@ 4.2 Picking Key Experts from Domain‑Specialized Experts
Concretely, taking math tasks as an example, we use the domain‑specialized experts identified in the previous analysis as candidate experts in each layer. We then run the model on 1000 randomly sampled instances from the MathQA (amini-etal-2019-mathqa, ) dataset. For each sample, we record the output logits of the original model, and then repeat the inference while pruning one candidate expert at a time. The perturbation of the output distribution is quantified using the KL divergence between the probability distributions of the Top‑1000 tokens from the original and pruned models (see Appendix B for implementation details of Top-1000 KL divergence).
@@
To test this hypothesis, we construct a failure set consisting of the Math‑500 (lightman2023lets, ) problems that the base model originally answers incorrectly (24 in total). For each candidate expert, we forcibly include it during routing and then re‑evaluate accuracy on this failure set. As shown in Figure 4 (right vertical axis), only those experts that induce substantial distribution shifts also yield meaningful accuracy gains when enhanced, whereas others contribute little to no improvement. In particular, expert L9E18, which caused the largest divergence, also increases the number of correct answers by 5 out of 24 when included. This provides strong evidence of the close correspondence between distribution shifts and functional importance, and offers a principled way to distinguish key experts. from the many domain‑specialized experts.
@@ 4.3 Enhancing Key Experts
Figure 3: Impact of pruning (measured by KL divergence, left axis) and enhancing (measured by accuracy gain on failure set of Math-500, right axis) for domain‑specialized experts.
Figure 4: Comparison of five designed enhancement methods for key experts, evaluated by accuracy on three widely used math benchmarks including AIME2024, AIME2025 and Math-500.
@@
A (Forced Addition): whenever a key expert is not selected, it is added alongside the chosen experts.
B (Forced Replacement): similar to A, but the key expert replaces the selected expert with the lowest routing weight.
C (Range-based Addition): if a key expert is not selected but appears within the router’s top‑$2k$ candidates, it is forcibly added.
D (Range-based Replacement): under the same condition as C, the key expert replaces the weakest selected expert.
E (Logit Bias): before routing, a fixed bias is added to the routing logits of key experts. In our implementation, the bias value is set to $20\%$ of the average weight among the top-$k$ selected experts.
We evaluate these methods on AIME2024, AIME2025, and Math‑500. For the two AIME datasets, results are averaged over five runs to mitigate small‑sample variance. Figure 4 summarizes the outcomes. Overall, strategy A brings consistent but modest gains. Strategy B, which forces replacement regardless of probability, is unstable and even reduces accuracy on AIME2024 and Math‑500, indicating that overly aggressive intervention can be harmful. By contrast, strategies C and D—enhancing only when the key expert is already close to the router’s top candidates—deliver the strongest improvements across all benchmarks. The difference between C and D is minor, but since D avoids expanding the number of active experts, it is preferable in practice. Strategy E shows less stable performance, as the bias magnitude is sensitive and difficult to tune.
@@ 5 Ban
Our dynamic expert pruning method is motivated by two key observations: 1) Different layers in an MoE model exhibit highly diverse sensitivity to expert pruning and thus should be treated differently. 2) During reasoning, the token-wise routing distributions also vary significantly: when most of the routing weight is concentrated on a few top experts, pruning has little effect; in contrast, when the weights are more evenly distributed, pruning becomes more harmful.
@@
We further analyze token‑level pruning sensitivity in long reasoning processes by examining how routing weights are distributed across experts. To quantify this, we compute the ratio between the cumulative routing weights of the top‑3 and top‑8 experts, averaged across all layers. Figure 5b plots these ratios for the first 200 generated tokens in a reasoning process from one GPQA‑Diamond sample. The results reveal substantial variability across tokens: some tokens concentrate most of their weights on a few experts and are thus robust to pruning, whereas others distribute weights more evenly and are therefore more sensitive.
@@
$S_{i,l}=\beta\cdot L^{\prime}_{l}+(1-\beta)\cdot T^{\prime}_{i},$
where $\beta\in[0,1]$ balances the contributions of layer- and token-level sensitivities; we set $\beta=0.5$ in our experiments to achieve a balanced trade-off. The hyperparameter $\lambda$ scales the combined sensitivity score and thus directly controls the aggressiveness of pruning. The dynamic number of selected experts is determined as
$S^{\prime}_{i,l}=\lambda\cdot S_{i,l},\qquad K_{i,l}=\left\lfloor K_{\min}+(K_{\text{base}}-K_{\min})\cdot S^{\prime}_{i,l}\right\rceil,$
where $K_{\text{base}}=8$ is the default number of experts and $K_{\min}=3$ is the minimum value. The choice of $K_{\min}$ is based on empirical inspection: reducing the number of selected experts below this threshold causes the model to collapse entirely. A detailed analysis of the choices of the hyperparameters $\lambda$ and $K_{\min}$ is provided in the LABEL:{sec:pruning_param}.

### v2

# Ban&Pick: Ehancing Performance and Efficiency of MoE-LLMs via Smarter Routing

@@ Abstract
Sparse Mixture-of-Experts (MoE) has become a key architecture for scaling large language models (LLMs) efficiently. Recent fine‑grained MoE designs introduce hundreds of experts per layer, with multiple experts activated per token, enabling stronger specialization. However, during pre‑training, routers are optimized mainly for stability and robustness: they converge prematurely and enforce balanced usage, limiting the full potential of model performance and efficiency at inference. In this work, we uncover two overlooked issues: (i) a few highly influential experts are underutilized due to premature and balanced routing decisions; and (ii) enforcing a fixed number of active experts per token introduces substantial redundancy. Instead of retraining models or redesigning MoE architectures, we introduce Ban&Pick, a post-training, plug-and-play strategy for smarter routing. Pick discovers and reinforces key experts—a small group with outsized impact on performance—leading to notable accuracy gains across domains. Ban further dynamically prunes redundant experts based on layer and token sensitivity, delivering faster inference with minimal accuracy loss. Experiments on fine-grained MoE-LLMs (DeepSeek, Qwen3) across math, code, and general reasoning benchmarks demonstrate that Ban&Pick delivers free performance gains and inference acceleration without retraining or architectural changes. For instance, on Qwen3-30B-A3B, it improves accuracy from 80.67 to 84.66 on AIME2024 and from 65.66 to 68.18 on GPQA-Diamond, while accelerating inference by 1.25× under the vLLM.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress (DBLP:conf/iclr/ZhouWLSLQLJSZ024, ), yet the ever-growing demand for higher capacity presents significant challenges for efficiency. As a promising solution, Mixture-of-Experts (MoE) architectures scale parameter count without proportionally increasing computation, by activating a small subset of experts for each input (artetxe-etal-2022-efficient, ).
@@
However, we argue that a critical but overlooked phenomenon exists: the routing strategies learned during pre-training cannot fully exploit this specialization. Specifically, we empirically discover that current routing strategies inadvertently limit the potential to leverage the most impactful experts while simultaneously introducing redundancy where many selected experts contribute minimally to the final output. Instead of attempting to redesign training objectives or expert architectures, we take a post-training perspective, aiming to unlock this untapped potential efficiently and effectively.
Delving into this question, we surprisingly find that within the most frequently selected experts of fine‑grained MoE models, a small fraction plays an outsized role: simply forcing these experts to be chosen for every input can noticeably boost accuracy on certain datasets. Motivated by this striking observation, we conduct an empirical analysis of expert usage across domains and discover that each domain consistently relies on its own set of frequently selected experts, which we refer to as domain‑specialized experts. Yet, a closer examination further shows that only a tight subset of these experts has a decisive impact on the final logits distribution, while the influence of the others remains marginal. We term these most critical ones key experts. Our Pick module is designed to reinforce their influence during routing, getting performance gains by emphasizing these key experts.
@@
Attempting to bridge the gap between routing strategies learned during pre-training—which prioritizes stability and balance—and optimal inference-time expert utilization, Pick and Ban empower truly impactful experts to contribute more while reducing participation from low-contribution experts:
  • Pick amplifies excellence by identifying experts that demonstrate superior performance on specific domains and strategically increasing their influence during inference, maximizing effectiveness.
  • Ban eliminates inefficiency by dynamically excluding experts that contribute minimally during inference, reducing computational costs while preserving essential capabilities.
Together, we introduce Ban&Pick, a unified framework that optimizes expert utilization in inference to improve both the accuracy and efficiency of fine-grained MoE without retraining or architectural changes. Evaluated on five widely used datasets spanning math, code, and general reasoning, the results demonstrate that on the Qwen3 series models, applying Pick alone brings an average 2.83% performance improvement. By further incorporating Ban, the joint Ban&Pick strategy achieves an average 1.41% improvement while simultaneously accelerating inference by 1.25× under the vLLM.
@@ 2 Related Works
Enhancing Expert Utilization in MoE.
Most existing work attempts to improve expert utilization efficiency in MoE models by redesigning training objectives or expert architectures during pre-training (yan2025tcmoe, ; huang-etal-2024-harder, ; jin2024moeacceleratingmixtureofexpertsmethods, ). However, while these approaches are insightful, they often sacrifice training stability for flexibility. Recently, (wang2025expertsneedsteeringthinking, ) proposes a post-training method designed for MoE models that employ the <think> token, enhancing reasoning by amplifying the weights of Cognitive Experts.
Dynamic Expert Pruning of MoE.
Dynamic expert pruning has emerged as a key technique for accelerating MoE inference. (lu-etal-2024-experts, ) proposed a pruning strategy that simply skips experts with less weights. Building on this, (huang2024mc, ) introduced a critical token protection mechanism. However, the substantial sensitivity differences across layers of MoE models and across tokens in long chain‑of‑thought reasoning are often overlooked, leaving room for further optimization.
3 Empirical Findings on Routing Limitations in Fine-grained MoE
Figure 1: The intuition and empirical findings on expert utilization. Left (Pick): forcibly activating key experts (e.g., E43) improves accuracy on Winogrande (sakaguchi2021winogrande, ). Right (Ban): reducing experts from 6 to 5 cuts 16.7% computation without accuracy loss.
3.1 Insights from Pre-training Patterns
Recent evidence from OLMoE (muennighoff2025olmoeopenmixtureofexpertslanguage, ) and OpenMoE (xue2024openmoe, ) suggests that during pre-training, routers converge much earlier than the experts. In particular, OLMoE reports that after only 1% of pre-training, around 60% of the routing decisions have already saturated, while OpenMoE finds that by 40% of pre-training, about 80% of the routing has stabilized. Such early fixation implies that routing decisions are largely determined before the experts have sufficiently matured. Moreover, the widely used balancing loss (lepikhin2020gshardscalinggiantmodels, ), while effective at preventing collapse to just a few experts, may also discourage tokens from concentrating on the most impactful experts. This implies that the current pre-training paradigm for MoE makes a trade-off: in exchange for training stability, it locks routers into decisions before experts have fully matured. Combined with balancing constraints, this underutilizes high-potential experts and inflats the role of those that contribute little at inference. This presents an opportunity from a post-training perspective:
@@ 3.1 Insights from Pre-training Patterns
3.2 Expert Utilization Inefficiencies and Optimization Potential
To validate the inefficiencies in expert untilization and better understand the optimization potential, we conducted two simple yet revealing experiments with the Deepseek‑V2‑Lite‑Chat model (activates 6 experts out of 64). Figure 1 demonstrates how we adjust expert utilization and its effects.
(1) Underutilization of the most impactful experts: we examined expert selection frequency on the C4 (dodge2021documentinglargewebtextcorpora, ) calibration set, identified the top-3 most frequently selected experts in each layer, and then conducted a grid search: each of these experts was manually added in addition to the original routing (E43 in Figure 1), and the resulting accuracy on Winogrande benchmark was recorded. The left part highlights a few experts contribute disproportionately: enhancing Layer‑1 Expert‑43 improved accuracy by +1.81%, L10E33 by +1.65%, and L25E11 by +1.26%. When these three were jointly enhanced, accuracy increased by +2.72% (71.67% $\rightarrow$ 74.39%), while most other high-frequency experts had negligible or noisy effects, indicating that certain impactful experts are insufficiently leveraged in the original routing strategy. We refer to these three experts as key experts.
(2) Computational waste from low-contribution experts: we directly reduced the number of selected experts per token from 6 to 5 and measured the resulting accuracy. Remarkably, as shown in the right side of Figure 1, this pruning reduced computation by 16.7% and surprisingly preserved full accuracy on Winogrande (71.67%), demonstrating redundancy in the original expert utilization.
Together, these experiments validate the inefficiencies in expert utilization predicted by pre-training analysis, and support the intuition of our proposed methods—Pick and Ban, respectively, which aim to optimize expert selection beyond pre-training constraints for better inference-time performance. This empirical study motivates us to design a unified framework that can pick key experts while banning redundant ones, aiming at improving both accuracy and efficiency.
@@ 4 Pick
In this section, we begin with an analysis of expert specialization in fine‑grained MoE models to verify the existence of domain-specialized experts, which forms the foundation for key experts to effectively function. Building on this analysis, we then introduce our Pick method in two steps: (i) identifying key experts from the set of domain‑specialized experts, and (ii) designing strategies to enhance these key experts so that their impact on model performance can be effectively amplified.
@@ 4.1 Expert Specialization in Fine-Grained MoE
First, we take a closer look at the degree of expert specialization in fine‑grained MoE models. Specifically, we analyze the Qwen3‑30B‑A3B model and record expert usage frequencies on task-related benchmarks: math tasks (GSM8K + Math-500), code tasks (MBPP (austin2021program, ) + Livecodebench), and a general task (GPQA-Diamond). The statistics cover all tokens in both the prefill and decode stages. For each type of task, we plot the per‑layer expert usage frequency and show 3 representative layers in Figure 2 (complete results in Appendix A).
Our analysis reveals that in each layer there exist experts whose usage frequency is significantly higher than average–often up to 6 times higher. Moreover, in most layers these high‑frequency experts vary across tasks. For instance, L16E95 is frequently activated only on math tasks, whereas L29E91 is almost exclusively activated on code tasks with an extremely high activation rate of 66.9%. In contrast, a small number of experts such as L4E48 remain highly active across all categories of tasks.
@@ 4.2 Picking Key Experts from Domain‑Specialized Experts
Concretely, taking math tasks as an example, we use the domain‑specialized experts identified in the previous analysis as candidate experts in each layer. We then run the model on 1000 randomly sampled instances from the MathQA (amini-etal-2019-mathqa, ) dataset. For each sample, we record the output logits of the original model, and then repeat the inference while pruning one candidate expert at a time. The perturbation of the output distribution is quantified using the KL divergence between the probability distributions of the Top‑1000 tokens from the original and pruned models (see Appendix B for implementation details).
@@
To test this hypothesis, we construct a failure set consisting of the Math‑500 (lightman2023lets, ) problems that the base model originally answers incorrectly (24 in total). For each candidate expert, we forcibly include it during routing and then re‑evaluate accuracy on this failure set. As shown in Figure 4 (right vertical axis), only those experts that induce substantial distribution shifts also yield meaningful accuracy gains when enhanced, whereas others contribute little to no improvement. In particular, expert L9E18, which caused the largest divergence, also increases the number of correct answers by 5 out of 24 when included. This provides strong evidence of the close correspondence between distribution shifts and functional importance, and offers a principled way to distinguish key experts from the many domain‑specialized experts.
@@ 4.3 Enhancing Key Experts
Figure 3: Impact of pruning (measured by KL divergence, left axis) and enhancing (by accuracy gain, right axis) for domain‑specialized experts.
Figure 4: Comparison of five designed enhancement methods for key experts, evaluated by accuracy on three widely used math benchmarks.
@@
  1. A.
    Forced Addition: Whenever a key expert is not selected, it is added alongside the chosen experts.
  2. B.
    Forced Replacement: Similar to forced addition, but the key expert replaces the selected expert with the lowest routing weight, maintaining the total number of selected experts.
  3. C.
    Range-based Addition: If a key expert is not selected (not in the top-$k$ candidates) but appears within the router’s top-$2k$ candidates, it is forcibly added.
  4. D.
    Range-based Replacement: Under the same condition as range-based addition, the key expert replaces the selected expert with the lowest routing weight.
  5. E.
    Logit Bias: Before routing, a fixed bias is added to the routing logits of key experts. In our implementation, the bias is set to $20\%$ of the average weight among the top-$k$ selected experts.
We evaluate these methods on AIME2024, AIME2025, and Math‑500. For the two AIME datasets, results are averaged over five runs to mitigate small‑sample variance. Figure 4 summarizes the outcomes. Overall, strategy A brings consistent but modest gains. Strategy B, which forces replacement regardless of probability, is unstable and even reduces accuracy on AIME2024 and Math‑500, indicating that overly aggressive intervention can be harmful. By contrast, strategies C and D—enhancing only when the key expert is already close to the router’s top candidates—deliver the strongest improvements across all benchmarks. The difference between C and D is minor, but since D avoids expanding the number of active experts, it is preferable in practice and used in our experiments. Strategy E shows less stable performance, as the bias magnitude is sensitive and difficult to tune.
@@ 5 Ban
Our dynamic expert pruning method is motivated by two key observations: 1) Different layers in an MoE model exhibit highly diverse sensitivity to expert pruning and thus should be treated differently. 2) During reasoning, the token-wise routing distributions also vary significantly: when most of the routing weight is concentrated on a few top experts, expert pruning causes minimal loss; in contrast, when the weights are more evenly distributed, expert pruning becomes more harmful.
@@
We further analyze token‑level pruning sensitivity in long reasoning processes by examining how routing weights are distributed across experts. To quantify this, we compute the ratio between the cumulative routing weights of the top‑3 and top‑8 experts, averaged across all layers. Figure 5b plots these ratios for the first 200 generated tokens in a reasoning process from one GPQA‑Diamond sample. The results reveal substantial variability across tokens: some tokens concentrate most of their weights on a few experts and are thus robust to expert pruning, whereas others distribute weights more evenly and are therefore more sensitive.
@@
$S_{i,l}=\lambda\cdot\tfrac{1}{2}\big(L^{\prime}_{l}+T^{\prime}_{i}\big),\quad\lambda\in(0,1)$
where we simply average layer- and token-level sensitivities to balance their contributions, and use $\lambda$ to control the aggressiveness of pruning. The dynamic number of selected experts is determined as:
$K_{i,l}=\left\lfloor K_{\min}+(K_{\text{base}}-K_{\min})\cdot S_{i,l}\right\rceil.$
Here $K_{\text{base}}=8$ is the default number of experts, and $K_{\min}=3$ serves as a safeguard to prevent selecting too few experts: from empirical inspection, we found that reducing the number of selected experts below this threshold leads to a complete collapse of the model. Details of the experimental validation for the setting of $K_{\min}$ and the choice of $\lambda$ are provided in the Appendix C.